OpenStudy (anonymous):

The following table represents the total cost, in dollars (y) to join a gym for x number of months.The cost includes a one time joining fee of $10 . Does the data in the table represent a proportional relationship or a non-proportional relationship? how do you know? x 1 2 3 4 5 y 25 40 55 70 85

OpenStudy (texaschic101):

make a proportion using the points and see if it is equal 1/25 = 2/40 --- cross multiply (25)(2) = (1)(40) 50 = 40 -- no...this is not proportional..it does not equal another example : x 1,2,3,4 y 3,6,9,12 1/3 = 2/6 -- cross multiply (3)(2) = (1)(6) 6 = 6 -- correct...this would be proportional